A client with dehydration is receiving continous iv fluids.1Liter q 6 hr. how many ml will the client receive in 8 hr
Answer #1

Solution:

The patient is receiving continuous I.V. 1liter in 6 hours.

1liter = 1000ml.

So, In 6 hours the patient receives 1000ml

Therefore, In 8 hours the patient will receive= 8× 1000= 8000ml

8000 ÷ 6 = 1333.33 ml.

The client will receive 1333.33 ml in 8 hours.